OpenStudy (anonymous):

PLEASE HELP IM STUCK WITH THIS ONE AND A COUPLE MORE What is the value of b? 11+4b=8b-13

OpenStudy (anonymous):

11+4b=8b-13 -4b -4b 11=4b-13 first you write it out like this do you understand r want me to explain it more?

OpenStudy (anonymous):

11+4b=8b-13 4b-8b=-13-11 -4b=-24 b=-24/-4 b=6

OpenStudy (anonymous):

@94634 so the anwer is 6?

OpenStudy (anonymous):

yes it is that's the correct working for it.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

no worries the answer is correct, u understand how i got it